EduFarming
Your list · 0हिं

Type at least two letters. Press Esc to close.

↑↓ navigate↵ openTAB cycle filter

Binomial & Poisson Distributions · Question 65 of 68

Question 65

HardOriginal question

XX follows a Poisson distribution with λ=1\lambda = 1. What is P(X≥2)P(X \ge 2)? (Take e−1=0.3679e^{-1} = 0.3679.)

Options for question 65

Show the answer and explanation
CorrectNot quiteAnswer C

Correct answer: C

Explanation

  1. P(X≥2)=1−P(X=0)−P(X=1)P(X \ge 2) = 1 - P(X = 0) - P(X = 1), and with λ=1\lambda = 1 both are e−1=0.3679e^{-1} = 0.3679.

  2. =1−0.7358=0.2642= 1 - 0.7358 = 0.2642.

Worth noting

0.6321 takes away only P(X=0)P(X = 0); that is P(X≥1)P(X \ge 1).

Concept
At least r events
Typical time
120 s
Your time
—
The rule this testsThe Poisson distribution
Practise this chapter →

Ten questions from Binomial & Poisson Distributions, at your own pace.

Practise this chapter →

Ten questions from Binomial & Poisson Distributions, at your own pace.